The lenghth of a rectangle garden is 9 ft less than 4 times its width. If the perimeter of the garden is 52 ft, find the dimensions of the garden
Answer by sulekha_hi(1)

let the width of the garden be x ft
length of the garden = 4x-9 ft
perimeter = 52 ft
acording to question,
perimeter of the rectangular garden = 2(length+breadth)
2(4x-9 + x)= 52
2 (5x-9)= 52
5x-9 = 26
5x = 26+9
5x=35
x=7
Width of the garden= 7 ft
length = 4* 7-9
= 19 ft